OpenStudy (micheal1234):

1.is (0,3) a solution to the equation y=x+3? yes or no 2.is (1,-4) a solution to the equation y=-2x yes or no 3.Look at the following points. (4,0), (3,-1), (6,3), (2,4) Which are solutions to y=x-4? choose all correct answers there are two correct answers. A. (6,3) B. (is (4,0) C. (3,-1) D. (2,-4) 4.There are 8 crackers in 1 serving, 16 crackers in 2 servings, 24 crackers in 3 servings, and so on. How many crackers are in a box of 12 servings? A.80 B.96 C.108 D.120

OpenStudy (anonymous):

to complete the problems 1 and 2, just substitute the first number in brackets into x and second into y. Solve and check if it is true or not.

OpenStudy (anonymous):

the same is for 3rd problem. Substitute instead of x and y each pair of numbers.

OpenStudy (micheal1234):

what do you mean i am not good with big words

OpenStudy (anonymous):

1. you have equation y=x+3. You have 0 for x and 3 for y. Substitute: 3 = 0 + 3 This is true.

OpenStudy (anonymous):

so (0, 3) is solution to equation y=x+3

OpenStudy (anonymous):

do the same thing with 2nd.